OpenStudy (anonymous):

Which property is shown by -5 x (6+7)= -5 x 6 +(-5) x 7 A: Commutative property of Addition B: Identity property of Addition C: Distributive Property D: Associative Property Of Addition

OpenStudy (zzr0ck3r):

The commutative property of addition says a+b=b+a The identity property of addition says a+0=a The distributive property says a(b+c)=ab+ac Associative property of addition says a+(b+c)=(a+b)+c

OpenStudy (anonymous):

I got the answer it is C